# Optimization of Excitation Waveforms for Maximum Instantaneous Field Intensity

### v2

@@ Abstract
This paper introduces a computational approach to identify performance constraints in the time-domain based on optimizing the excitation waveform. The method builds on an optimization algorithm that has been employed for decades to establish fundamental limits in the frequency domain and this paper showcases its first comprehensive application to time-domain pulses. The method is applied to arbitrarily polarized multiport antennas and arrays. The demonstration performed is based on finding an antenna’s maximum peak radiation intensity in a given direction and time with limited total input energy available. To highlight the generality of the approach, an analysis on finding optimal illumination for antiferromagnetic memory switching is conducted.
@@
Optimal control, excitation, pulse radiation, peak power, electromagnetic radiation, antenna arrays, magnetic memories.
@@ I Introduction
Short EM (EM) pulse radiation is of significant importance in many fields, including applications involving impulse, short-pulse, or non-sinusoidal radars [1, 2, 3], ground penetrating radars [4] or through-wall imaging radars [5]. In addition to radar applications, an important discipline within the field is EM immunity testing [6]. Moreover, it has played a crucial role in communication systems that seek to increase information transmission rates [1], EM pulse applications, broadband arrays and high-energy directed beam systems [7, 8, 9, 10]. Apart from radioengineering, pulses are used across the whole spectrum of physics including antiferromagnetic switching which is analyzed in this manuscript.
Traditionally designed antennas using continuous wave analysis and synthesis techniques are frequently used to radiate short-duration pulses. Antennas with wide bandwidth are typcially used, relying on the belief that they ensure sufficient performance. A more advanced approach involves matching an antenna excitation signal with its dispersion characteristics. Such an approach yields significantly better results [1, 11, 7]. There are no widely recognized standards for evaluating the performance of an antenna that radiates pulses, although the following are generally accepted: the pulse distortion in the radiation should be minor, the energy of the reflected signal at the input should be small, or the amplitude of the radiated pulse should be large [4].
This paper establishes a computational approach to determine optimality in the radiation of directional pulses, generalizes the methodology, and allows the tool to be used in unconventional scenarios which had not been previously possible. The tool can assess the quality of an arbitrary antenna and give its optimal excitation.
@@
While co-design approaches that simultaneously optimize waveform and antenna design may approach globally optimal solutions, this work focuses specifically on waveform optimization for fixed antenna designs. By decoupling waveform optimization from antenna design, the method ensures broad applicability and low computational complexity.
To find the best possible excitation waveform, the optimization framework stems from the development of fundamental bounds on antenna performance in the frequency domain. By including commonly used computation methods, modern approaches, and evaluation techniques such as the EFIE (EFIE[12], MoM (MoM[13], convex optimization [14] and powerful computation units, limitations are often found on arbitrary quantities in various scenarios. A representative example is when efficiency, Q factor, and gain [15, 16] are intrinsically connected to directional pulse radiation which natively requires antennas with sufficient performance in these quantities. The method was initially introduced as antenna current optimization [17], and is based on current density being expressed in terms of EFIE and MoM representing dof (dof) in convex optimization. In state of the art in fundamental bounds in electromagnetism [18], quantity trade-offs [15, 19], tighter bounds including additional constraints [19], or quantities in specific scenarios, such as implanted antennas [20], are evaluated. The phenomenon is also well developed in fields unrelated to antennas, i.e., cloaking [21], heat transfer [22], magnetic trapping [23], and harmonic generation [24].
A few works addressing optimal excitation in the time domain have been done, notably those considering radiated pulse optimization in [8, 25], where up-to-date methodologies are not used. Initial attempts to use modern tools in the optimization of radiated pulses are presented in [26]. The results show optimality in peak radiation intensity in a given direction of linearly polarized dipole-type antennas, such as the dipole, bowtie, and exponentially tapered bowtie. Evaluation consists of a large number of frequency samples computed using EFIE and MoM to optimize the excitation using the  quadratically constrained quadratic program (QCQP[14, 27] in the frequency domain and then mapping by inverse fast Fourier transform (iFFT) to the time domain [26]. The extension of the method to arbitrarily polarized antennas was initialized in [28].
In this paper, in addition to the applicability of the approach to an almost arbitrary problem using any simulation software, multiport antennas and arrays are the novel features described, as discussed in Section IV. The section III presents the framework in detail and follows the flow chart in Fig. 1 where individual steps are given. Most steps are performed using well-developed, proven techniques in the frequency domain for antenna analysis and computation of fundamental bound. The procedure allows for direct control of the frequency range of the optimal excitation pulse which is always band limited, i.e., angular velocity $\omega=2\pi f$ and frequency $f$ are considered from $f_{\mathrm{min}}$ to $f_{\mathrm{max}}$, given by realistic values [^1] For instance, regarding antenna radiation, the frequency band is constrained using step-recovery diodes [29] with $f_{\mathrm{min}}=0\,\mathrm{Hz}$ and $f_{\mathrm{max}}=3.8\,\mathrm{GHz}$. for pulse generators.
antenna S.1 EM solver characte-ristics S.2 interpolation operators $(\omega)$ S.3 projection basis functions S.4 optimization QCQP solution $(\omega)$ S.5 iFT solution $(t)$
Figure 1: Flow chart of the computation steps S.1 – S.5, starting from the given antenna and port with the characteristic impedance of the feed transmission line and finishing with the optimal solution in the time domain, i.e., the optimal excitation and performance value.
@@
The remainder of the paper is organized as follows. Section II describes the history of how to generate excitation for optimal pulses. The established computation framework for optimal directionally radiated pulses is presented, tested, and verified in an example in Section III, whereby a comparison of the performance of different antenna designs allows the best one to be chosen, and assists in deciding on the optimal layout of an array for pulsed radiation. Using the same methodology, the optimal illumination for antiferromagnetic memory switching is found in Section IV. Capabilities, limitations, and possible extensions are discussed in Section V.
Although this work primarily focuses on optimizing excitation waveforms to achieve power metrics in far-field scenarios. The proposed framework is easily adjustable for near-field energy focusing or similar applications where precise waveform shaping is crucial, given the knowledge of the impulse response in the respective region. However, this work does not address the practical aspects of designing pulsed systems but instead provides a versatile theoretical methodology that is applicable in diverse contexts.
@@ II Historical Background
The critical advancement to uncovering the optimality of antenna pulse radiation lay in using an excitation signal precisely tailored to the specific dispersion characteristics of the antenna [1]. Achieving this would result in a flat frequency response producing the highest possible pulse amplitude and maximizing the strength of the electric field at a given moment and point in the far field. For arbitrarily shaped antennas, with a particular emphasis on planar pulse radiators, the search for optimality is detailed in [1] and provides insight into performance maximization related to size and bandwidth. The methodology assumes an arbitrary planar current distribution confined to the design region. Although this understanding of bounds provides valuable insight, it falls short of synthesizing practical and applicable designs.
@@
In [7], a distinct performance maximum of pulse radiation is initially introduced to examine the optimal input voltage. The metric and computational approach studied remain consistent. This exploits the computation in the frequency domain, using iFFT, and applies variational optimization to maximize the amplitude of the radiated electric field at a specified time and position in the far field with constrained energy [^2] Constraining the root mean square current amplitude over the surface of the radiator was implemented in [1, 7]. and the bandwidth of the applied signal. Compared to [1], the current in [7] is physically realizable bringing this solution closer to practical implementation as long as the feed voltage can be generated or sufficiently approximated. Although the solution in [7] refers to a dipole antenna, it accommodates arbitrary feed point locations, lengths, or impedance loadings. In these works, the current distribution is determined using EFIE and the Galerkin method to employ piecewise sinusoidal expansion [30].
@@
The progression of the study in determining the optimal transient radiation of arbitrary antennas is documented in [8, 31] and extends the analysis to three-dimensional antennas compared to planar radiators in [1]. The objective metric remains the same as before. Furthermore, the study addresses the maximum possible radiated energy density within a specified time interval while considering the effects of introducing constraints on the sidelobes. Physical realizability is emphasized by limiting Q or super-directivity effects. Fundamental limitations are juxtaposed with the performance of optimally fed antennas [7, 31] and dipole arrays [9], all of the same size [8, 31]. This work builds on [32] and employs spherical mode expansion. Limiting the number of spherical mode expansion terms ensures a low antenna Q and mitigates impractical supergain issues. The Chebyshev synthesis procedure is used to constrain the energy radiated in the side lobes.
@@
Supplementary to [7] and in alignment with [8], optimal excitations for dipole arrays aimed at maximizing the amplitude of the transient radiated electric field at a specified time and position in the far field are detailed in [9], and maximizing energy density in a given time interval at a fixed far-field position are detailed in [2]. As with [7], the MoM is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling. Piecewise sinusoidal expansion modes are used, reflecting the approach in [7]. The relationship between the total radiated field and terminal voltages is established by superposing the far fields of each terminal. This analysis uses a thin-wire approximation for the dipoles and delta gaps to model their feed-voltage sources. The dof in this study encompass the magnitudes and phases of the terminal voltages. Furthermore, the investigation is extended to the synthesis of optimal sidelobe patterns.
Alternatively, one might consider whether the radiated electric field at a far zone is equal to a predetermined function of time rather than focusing on maximizing field intensity. This question is addressed in [10]. In specific applications, such a criterion may be more suitable than maximizing the radiated field amplitude or energy density at a particular time and position in the far zone. The solution to this problem is carried out entirely in the time domain by minimizing a particular norm of the difference between the required and actual electric fields at a point lying in the far zone. The paper tackles the inverse source problem, specifically constructing the input voltage for a given far-zone electric field function of time via an iterative optimization algorithm based on the steepest descent.
@@
Given that all previous work has focused on radiation in the far-field zone, the study of the near-field zone remains open. Concentrating EM energy in small regions for short intervals has emerging applications in biomedical applications, such as hyperthermia therapy to treat cancer and other diseases [3]. The approach to achieve a high concentration of EM energy in small regions involving free space and non-dispersive materials is presented in [3]. The framework is demonstrated with linear arrays. Intuitively, the timing of the sources must be adjusted to compensate for differences in the propagation path lengths so that transmitted pulses from the array elements are coherently added at the target location.
@@
Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas, in contrast to the work mentioned above. In ultra-wideband (UWB) radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [35]. Unbeatable target metrics are achieved by finding the best-performing waveform considering the constraints on energy in different stages of the channel. General optimization results are derived for arbitrary antennas, considering the effects of the generator and load impedances using variational methods. The approach is demonstrated with dipoles using MoM [^3] Piecewise sinusoidal basis is used as well, as in [7, 9]. solutions providing the necessary transfer functions and input impedances from the EFIE analysis with the study enriched by closed-form results for short-dipole antennas. Ideally, an UWB transmitter’s radiated power spectral density should be as flat as possible. However, in reality, UWB systems are limited by the performance of the antennas which is far from ideal [35].
The precise metric for the sharpness of the received antenna voltage amplitude and its optimization are detailed in [36]. The objective is the maximal derivative of the receiving voltage waveform. A substantial improvement is observed compared to [35] with results validated through measurement.
III Framework
@@
This section develops and verifies the general computational approach for optimal excitation and showcases it on particular example of pulsed radiation in a chosen far-field direction and time. Various goals and objective functions with multiple constraints can be considered within the approach presented in this article. The essential user input for the computation is the relation between the excitation and the target metric, i.e., a solution to Maxwell’s equations. Knowing the system response, almost arbitrary linear scenarios can be studied, including optimization of radiated field intensity, power delivery efficiency, or waveform shaping tailored to specific spatial or temporal requirements. This versatility makes the framework applicable to scenarios such as pulse radiation enhancement and electromagnetic field focusing in localized regions. In addition to the examples given in the manuscript, notable applications might be the wireless power delivery from the port of an antenna to the port of another antenna and the creation of hot spots in the near field. The practical usage of the method is given in Section IV where several applications are considered.
III-A Methodology
A representative goal used to demonstrate the methodology, which is widely analyzed in this manuscript, is to find an antenna’s maximum peak radiated power in a given direction and time with limited available total input energy available.
@@ III-A Methodology
$\begin{split}\max\limits_{\mathbf{a}(t)}\ &r^{2}\boldsymbol{r}_{0}\cdot\boldsymbol{S}(\boldsymbol{r},t_{0}+r/c),\\ \mathrm{s.t.}\ &W_{\mathrm{inc}}=W_{0},\\ &f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right],\end{split}$ (1)
where
$\boldsymbol{r}_{0}\cdot\boldsymbol{S}(\boldsymbol{r},t)=\dfrac{1}{Z_{0}}\left|\boldsymbol{E}_{\infty}(\boldsymbol{r},t)\right|^{2}.$ (2)
The incident waves $\mathbf{a}(t)$ on the feeding transmission line connected to the antenna ports are optimized to maximize radial component $\boldsymbol{r}_{0}$ of Poynting vector $\boldsymbol{S}$ at given point $\boldsymbol{r}$ and time $t_{0}+r/c$ which is normalized by the squared distance from source $r^{2}$. For convenience, the observation time is shifted by the free-space propagation delay caused by the speed of light $c$. Impedance $Z_{0}$ in (2) denotes the impedance of the free space and $\boldsymbol{E}_{\infty}$ denotes the electric intensity in the far-field region. Total incident wave energy $W_{\mathrm{inc}}$ is set to be total available energy $W_{0}$. This work considers $50$-Ohm feeding transmission lines and energy $W_{0}=10^{-10}\,\mathrm{J}$.
The solution to such a problem can be achieved by taking the following steps:
@@
    First, an antenna with defined ports and feeding transmission lines has to be chosen. The setup is analyzed in an arbitrary EM (EM) solver, which can compute the far field in the desired direction for a given frequency and incident waves on the feed lines to the antenna ports. The EM solver provides a set of far-field values. An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [37] (CST). This part of the process is computationally the most time-consuming. The computational demands grow with wider frequency ranges $f\in\left[f_{\mathrm{min}},f_{\mathrm{max}}\right]$, but these, on the other hand, introduce more dof in the waveform design providing greater flexibility in shaping the radiated pulses.
    The electric field in the far-field zone expressed in the frequency domain reads
    $\boldsymbol{E}_{\infty}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega)\mathrm{e}^{-\mathrm{j}\omega r/c},$ (3)
    where $\mathbf{H}$ is a far-field transfer matrix obtained from the chosen solver and $a(\omega)$ represents incident waves in the frequency domain (for example, power waves in the feeding transmission lines), i.e., Fourier transform of $a(t)$. An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A. Apart from the quantities needed for optimization, antenna properties such as the reflection coefficient or gain at given frequencies can be examined to provide insight into the spectral selectivity of the antenna.
@@
    The far-field data are interpolated using rational fitting (sparse data) or linear interpolation (dense data) to provide a smooth function of frequency. In this step, the operator for total incident energy is prepared, namely the identity operator in the quadratic form of the incident waves meaning the incident energy is written as a spectral integral
    $W_{\mathrm{inc}}=\dfrac{1}{4\pi}\int\limits_{-\infty}^{\infty}\mathbf{a}^{\mathrm{H}}\left(\omega\right)\mathbf{a}\left(\omega\right)\mathrm{d}\omega.$ (4)
@@
    The spectral functions representing the necessary operators for optimization, i.e., total incident energy, and the far field are projected onto a set of continuous-band-limited basis functions. The projection follows Galerkin method [38, 13]. The set of basis functions $\xi_{n}(\omega)$ is described in Appendix B. For all considered examples, number of basis functions [^4] Later, the number of basis functions is the number of dof in the optimization and determines the time needed for construction $\mathcal{O}(N^{2})$ of matrices and solution $\mathcal{O}(N^{3})$. $N\leq 120$ was sufficient. The projection results in expansion coefficients of the basis function grouped in column vector $\mathbf{q}$, where
@@
    The vector is constructed in segments with each segment comprising the expansion coefficients of the basis functions for each feed line $p$. Related matrices $\mathbf{W}$
    $W_{\mathrm{inc}}=\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}$ (6)
    (incident energy) and $\mathbf{S}$ (far-field power) are derived via Galerkin method from (4) and (2) together with (3), respectively.
    As an example, the incident energy (4) is represented by matrix $\mathbf{W}$, the components of which read
    $\mathbf{W}_{pm,qn}=\dfrac{\delta_{pq}}{4\pi}\int\limits_{-\infty}^{\infty}\xi_{m}^{*}\left(\omega\right)\xi_{n}\left(\omega\right)\mathrm{d}\omega,$ (7)
    where $pm,qn$ corresponds to the matrix rows and columns according to feeding lines $p,q$ and basis function $m,n$, respectively. Kronecker delta $\delta_{pq}$ is zero except when $p$ and $q$ refer to the same port. Since the ports are separated and the basis functions are orthogonal to each other with respect to the given integral, the identity matrix represents the total incident energy matrix $\mathbf{W}$. Equivalently projction of other operators considered in this methodological example are presented in Appendix C.
@@
    The expansion coefficients are optimized to maximize peak radiated power at the given time and direction according to (1) and the equivalent QCQP reads
    $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{S}\mathbf{q},\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{W}\mathbf{q}=W_{0}.\end{split}$ (8)
    The optimization task is resolved using standard methods of the QCQP. The optimization problem is convex, so the solution is unique [14], which ensures that the maximum is truly achieved [^5] The optimization problem is equivalent to that for determining performance limitations on antenna metrics in the frequency domain [19].. Within this paper, the optimization is more than a thousand times faster than the computation of the impulse response of the system. For more details about the convex optimization framework, see [39].
@@
    The solution is found as optimal vector $\mathbf{q}$ in the frequency domain. The linear combination on the feed line then gives the optimal feeding vector which results in maximum peak radiated intensity in the frequency domain using the operators of step S.2.
    The solution is transformed to the time domain via inverse Fourier transform (iFT) either as a single function or by weighted summation images of the basis functions $\xi_{n}(t)$. The far field is also transformed to depict the radiated pulse waveform in the time domain.
Although described for far-field radiation scenarios, the methodology explained above is general and can easily be applied to different directions and extended to applications such as near-field energy focusing, wireless power transfer, or UWB radars. However, we intentionally omit design concerns related to these systems, such as topology optimization of the radiating and receiving antennas. We focus solely on feeding waveform optimization, which is, unlike topology optimization, a convex optimization problem.
III-B Verification

### v3

# Joint Signal and Topology Optimization for Maximum Instantaneous Field Intensity

@@ Abstract
This paper introduces a computational approach to identify performance constraints in the time-domain, offering a way to design systems in pulse operation. This work presents a comprehensive application of convex optimization to determine fundamental bounds on time-domain waveforms. The approach is applied to arbitrarily polarized multiport antennas and arrays, demonstrating their capability in maximizing peak radiation intensity in a specified direction and time under energy constraints. This methodology allows us to consider matching, which is crucial in such applications. To highlight the generality of the approach, receiving systems are also studied on an example of finding optimal illumination for antiferromagnetic memory switching. Thanks to its efficacy, this work enables joint optimization of excitation and system parameters.
@@
Co-design, optimal control, excitation, antenna topology, pulse radiation, peak power, magnetic memories.
@@ I Introduction
Short EM (EM) pulse radiation is of significant importance in many fields, including applications involving impulse, short-pulse, or non-sinusoidal radars [1, 2, 3], ground penetrating radars [4] or through-wall imaging radars [5]. In addition to radar applications, an important discipline within the field is EM immunity testing [6]. Moreover, it has played a crucial role in communication systems that seek to increase information transmission rates [1], EM pulse applications, broadband arrays and high-energy directed beam systems [7, 8, 9, 10]. Apart from radioengineering and far-field operation, EM pulses are used across the whole spectrum of physics, including antiferromagnetic switching [11], which is analyzed in this manuscript.
Traditionally designed antennas using continuous wave analysis and synthesis techniques are frequently used to radiate short-duration pulses. Antennas with wide bandwidth are typically used, relying on the belief that they ensure sufficient performance. A more advanced approach involves matching an antenna excitation signal with its dispersion characteristics. Such an approach yields significantly better results [1, 12, 7]. There are no widely recognized standards for evaluating the performance of an antenna that radiates pulses, although the following are generally accepted: the pulse distortion in the radiation should be minor, the energy of the reflected signal at the input should be small, or the amplitude of the radiated pulse should be large [4].
This paper establishes a computational approach to determine optimality in the instantaneous field intensity, generalizes the methodology, and allows the tool to be used in unconventional scenarios which had not been previously possible. The tool can assess the quality of an arbitrary pulse system and give its optimal excitation. The approach is coupled with system design process to reach better performance.
@@
A few works addressing optimal excitation in the time domain have been done, notably those considering radiated pulse optimization in [8, 13], where up-to-date methodologies are not used. Initial attempts to use modern tools in the optimization of radiated pulses are presented in [14, 15, 16]. These efforts, as well as this manuscript, build upon earlier developments of fundamental bounds in the frequency domain [17, 18], which established a foundation for waveform optimization using convex optimization techniques [19].
@@
This paper demonstrates that the proposed method is adaptable to a variety of problems involving one or more sources using different simulation software. It also acts as a basis for automated co-design of pulse excitation and system topology. The method and examples in the manuscript are organized as follows. Section II describes the history of how to generate excitation for optimal pulses. The established computation framework is presented in Section III, and tested on an example of optimal directionally radiated pulses in Section IV. A comparison of the performance of different antenna designs is shown in Section V. Using the same methodology, the optimal illumination for antiferromagnetic memory switching is found later in Section VI illustrating the usage of the methodology for the nearfield and receiving parts of a system. The co-design, including pulse and topology optimization, is presented on an antenna array with multiple ports, demonstrating the broad applicability of the approach later in Section VII. Capabilities, limitations, and possible extensions are discussed in Section VIII.
@@ II Historical Background
The critical advancement to uncovering the optimality of antenna pulse radiation lies in using an excitation signal precisely tailored to the specific dispersion characteristics of the antenna [1]. Achieving this would result in a flat frequency response producing the highest possible pulse amplitude and maximizing the strength of the electric field at a given moment and point in the far field. For arbitrarily shaped antennas, with a particular emphasis on planar pulse radiators, the search for optimality is detailed in [1] and provides insight into performance maximization related to size and bandwidth. The methodology assumes an arbitrary planar current distribution confined to the design region. Although this understanding of bounds provides valuable insight, it falls short of synthesizing practical and applicable designs.
@@
In [7], a distinct performance maximum of pulse radiation is initially introduced to examine the optimal input voltage. The metric and computational approach remain consistent. This exploits the computation in the frequency domain, using inverse fast Fourier transform (iFFT), and applies variational optimization to maximize the amplitude of the radiated electric field at a specified time and position in the far field with constrained energy [^1] Constraining the root mean square current amplitude over the surface of the radiator was implemented in [1, 7]. and the bandwidth of the applied signal. Compared to [1], the current in [7] is physically realizable, bringing this solution closer to practical implementation as long as the feed voltage can be generated or sufficiently approximated. Although the solution in [7] refers to a dipole antenna, it accommodates arbitrary feed point locations, lengths, or impedance loadings. In these works, the current distribution is determined using EFIE (EFIE) and the Galerkin method to employ piecewise sinusoidal expansion [20].
@@
The progression of the study in determining the optimal transient radiation of arbitrary antennas is documented in [8, 21] and extends the analysis to three-dimensional antennas compared to planar radiators in [1]. The objective metric remains the same as before. Furthermore, the study addresses the maximum possible radiated energy density within a specified time interval while considering the effects of introducing constraints on the sidelobes. Physical realizability is emphasized by limiting Q or super-directivity effects. Fundamental limitations are juxtaposed with the performance of optimally fed antennas [7, 21] and dipole arrays [9], all of the same size [8, 21].
@@
Supplementary to [7] and in alignment with [8], optimal excitations for dipole arrays aimed at maximizing the amplitude of the transient radiated electric field at a specified time and position in the far field are detailed in [9], and maximizing energy density in a given time interval at a fixed far-field position are detailed in [2]. As with [7], the MoM (MoM) is employed to calculate the currents on all dipoles, taking into account the effects of mutual coupling.
@@
Given that all previous work has focused on radiation in the far-field zone, the study of the near-field zone remains open and can be addressed by the framework within this manuscript. Concentrating EM energy in small regions for short intervals has emerging applications in biomedical applications, such as hyperthermia therapy to treat cancer and other diseases [3].
Moving to communication systems, optimizing the transient waveform should consider both the transmitting and receiving antennas. In contrast to the work mentioned above, such an application can be covered by the approach proposed within this manuscript. In ultra-wideband (UWB) radio systems, the crucial metric is the maximum amplitude of the received antenna voltage, its sharpness, or energy, as addressed in [15].
@@
Unbeatable target metrics are achieved by finding the best-performing waveform considering the constraints on energy in different stages of the channel. General optimization results are derived for arbitrary antennas, considering the effects of the generator and load impedances using variational methods. The approach is demonstrated with dipoles using MoM [^2] Piecewise sinusoidal basis is used as well, as in [7, 9]. solutions providing the necessary transfer functions and input impedances from the EFIE analysis with the study enriched by closed-form results for short-dipole antennas. Ideally, an UWB transmitter’s radiated power spectral density should be as flat as possible. However, in reality, UWB systems are limited by the performance of the antennas which is far from ideal [15]. To achieve higher performance, antenna–pulse co-design is demanded.
III Waveform Optimization
@@
This section develops the general computational approach to maximize the instantaneous field intensity. Various goals and objective functions with multiple constraints can be considered within the approach presented in this article. The essential user input for the computation is the relation between the excitation and the target metric, i.e., a solution to Maxwell’s equations. Arbitrary linear scenarios can be studied, including optimization of radiated field intensity, power delivery efficiency, or waveform shaping tailored to specific spatial or temporal requirements. The methodology developed in this Section ensures that globally optimal time domain excitation is found for a given system.
The selected objective function $A(\mathbf{a})$ should be either linear or quadratic in the dof (dof), such as instantaneous power flux at given point, and is optimized under a series of constraints, $B_{i}(\mathbf{a})$, that are also quadratic or linear, notable examples being incident energy, delivered energy at a specific time frame, and field values at particular times and positions. The dof are represented by incident power waves [24] collected in vector $\mathbf{a}$. In contrast to previous work, this methodology allows us to consider matching, which, as shown in subsequent sections, is essential for high performance in the pulse regime.
@@
$\begin{split}\max\limits_{\mathbf{a}\left(t\right)}\ &A(\mathbf{a}),\\ \mathrm{s.t.}\ &B_{i}(\mathbf{a})\leq 0,\forall i\\ &\omega\in\left[\omega_{\mathrm{min}},\omega_{\mathrm{max}}\right].\end{split}$ (1)
The pulse optimization problem is addressed in the Fourier domain for $\mathbf{a}\left(\omega\right)$ by employing established techniques of convex optimization. The solution is further constrained by the available frequency band determined by the specific technology used in the design. Specifically, the angular frequency $\omega=2\pi f$ and the frequency $f$ are considered within the range from $\omega_{\mathrm{min}}$ to $\omega_{\mathrm{max}}$, based on practical values [^3] For example, in antenna radiation applications, the frequency band is limited using step-recovery diodes [25], where $f_{\mathrm{min}}=0\,\mathrm{Hz}$ and $f_{\mathrm{max}}=3.8\,\mathrm{GHz}$. for pulse generators.
The problem can be tackled by taking and performing the following steps which are also depicted in the flow chart in Fig. 1.
system S.1 EM solver characte-ristics S.2 interpolation operators $(\omega)$ S.3 projection basis functions S.4 optimization QCQP solution $(\omega)$ S.5 iFT solution $(t)$ S.6 redesign
Figure 1: Flow chart of the computation steps S.1 – S.6, starting from the given antenna and port with the characteristic impedance of the feed transmission line and finishing with the optimal solution in the time domain, i.e., the optimal excitation and performance value which may show a need of redesigning the system.
@@
    First, a system with defined excitation forms and input channels must be selected. The setup is then analyzed using an arbitrary EM solver capable of computing the resulting field intensity at a specified location for a given frequency and a set of incident power waves. The solver provides a set of field values corresponding to different excitations. An arbitrary method, capable of evaluating target quantities for a given excitation, can be used, i.e., finite differences in time domain (FDTD) in CST studio suite [26] (CST). This part of the process is computationally the most time-consuming. The computational demands grow with wider frequency ranges $\omega\in\left[\omega_{\mathrm{min}},\omega_{\mathrm{max}}\right]$, but these, on the other hand, introduce more dof in the waveform design providing greater flexibility in shaping the pulses.
    The provided field vector [^4] It is not strictly required to focus on fields; instead, parameters such as voltage at a port, outgoing power waves along a transmission line, or other related metrics can be analyzed. The crucial aspect is the linear relationship between these metrics and incident waves, as described by the transfer matrix. expressed in the frequency domain reads
    $\boldsymbol{E}(\boldsymbol{r},\omega)=\mathbf{H}(\boldsymbol{r},\omega)\,\mathbf{a}(\omega),$ (2)
    where $\mathbf{H}$ is a transfer matrix obtained from the chosen solver. An example of getting the transfer matrix is the combination of the EFIE and the MoM detailed in Appendix A. Apart from the quantities needed for optimization, system properties can be examined to provide insight into the spectral selectivity of the analyzed system.
@@
    The field data are interpolated using rational fitting (sparse data) or linear interpolation (dense data) to provide a smooth function of frequency. In this step, the operator for energy and field is prepared, the operators are listed in Appendix B.
@@
    The spectral functions representing the necessary operators for optimization, included in the objective function and constraint are projected onto a set of continuous-band-limited basis functions. The projection follows Galerkin method [27, 28]. The set of basis functions $\xi_{n}(\omega)$ is described in Appendix C. For all considered examples, number of basis functions [^5] Later, the number of basis functions is the number of dof in the optimization and determines the time needed for construction $\mathcal{O}(N^{2})$ of matrices and solution $\mathcal{O}(N^{3})$. $N\leq 120$ was sufficient and utilizing additional basis functions does not change the result. The projection results in expansion coefficients collected in column vector $\mathbf{q}$, where
@@
    The vector is constructed by blocks, each block representing one source. This allows to express functionals from Appendix B as quadratic forms in vectors $\mathbf{q}$. Taking energy as an example, the quadratic form reads
    $W(\boldsymbol{r},T)=\mathbf{q}^{\mathrm{H}}\mathbf{W}(\boldsymbol{r},T)\mathbf{q}$ (4)
    with matrix elements $\mathbf{W}_{pm,qn}$ where $pm,sn$ corresponds to the matrix rows and columns according to sources $p,q$ and basis function $m,n$, respectively.
@@
    The expansion coefficients are optimized to maximize the target performance metric according to (1) which is, using projections described in Step S.3, transformed into a solution to quadratically constrained quadratic program (QCQP) [^6] In later examples equality constraint is used, because the optimum is always reached with the equality. and reads
    $\begin{split}\max\limits_{\mathbf{q}}\ &\mathbf{q}^{\mathrm{H}}\mathbf{A}\mathbf{q}+\mathrm{Re}{\mathbf{a}_{i}^{\mathrm{H}}\mathbf{q}}+\alpha,\\ \mathrm{s.t.}\ &\mathbf{q}^{\mathrm{H}}\mathbf{B}_{i}\mathbf{q}+\mathrm{Re}{\mathbf{b}_{i}^{\mathrm{H}}\mathbf{q}}+\beta\leq 0,\forall i\end{split}$ (5)
    where one of the matrices $\mathbf{B}$ must be full rank and positive definite. This one is typically the total incident energy.
    The optimization task is resolved using standard methods of the QCQP. The optimization problem is convex, so the solution is unique [19], which ensures that the maximum is truly achieved [^7] The optimization problem is equivalent to that for determining performance limitations on antenna metrics in the frequency domain [29].. Within this paper, the optimization is more than a thousand times faster than the computation of the impulse response of the system (depending on discretization in spatial and frequency or domain). For more details about the convex optimization framework, see [30].
@@
    The solution to (5) is found as optimal vector $\mathbf{q}$. The desired optimal excitation is then found as
    $a_{p}(t)=\sum\limits_{n}q_{pn}\xi_{n}(t),$ (6)
    where inverse Fourier transforms of basis functions $\xi_{n}(\omega)$ are known.
  6. S.6
    If the performance of the current system with the optimal excitation is not sufficient, a redesign is needed, since solely by excitation, the performance cannot be enhanced further. For that, automated computer-powered system optimization can be employed. Automated co-design of the excitation and system topology then allows for even better performance.
    The automated co-design is computationally demanding and is realistic only when globally optimal excitations are found quickly. This is the case of a solution to QCQP developed in previous steps.
IV Pulse Optimization Verification